Tedco produces a teddy bear that sells for $24. Each teddy bear costs the company $8 to manufacture and there is an initial start-up cost of $7200.
If they want to make an annual pro t of $54 000, how many teddy bears do they need to sell?

Answer by ankor@dixie-net.com(22740)

Tedco produces a teddy bear that sells for $24. Each teddy bear costs the company $8 to manufacture and there is an initial start-up cost of $7200.
If they want to make an annual pro t of $54 000, how many teddy bears do they need to sell?
:
let t = no. of bears to accomplish this
Revenue - cost = profit
24t - (8t+7200) = 54000
24t - 8t - 7200 = 54000
16t = 54000+7200
16t = 61200
t = 61200/16
t = 3825 bears have to be made
:
:
Check on your calc
24(3825) - (8(3825)+7200) = 54000